OpenStudy (zachzach):

Let f(x) = twenty four over the quantity of 3 x minus two. Find f(−2). −6 −3 3 6

MARC (marc_d):

\[f(x)=\frac{24}{3x-2}\]

MARC (marc_d):

sub. x=-2 into the function\[f(\color{red}{-2})=\frac{24}{3(\color{red}{-2})-2}\]

MARC (marc_d):

\[f(-2)=\frac{24}{-6-2}\]

MARC (marc_d):

\[f(-2)=\frac{24}{-8}\]

MARC (marc_d):

simplify it \[f(-2)=\frac{3\times8}{-1\times8}\]

MARC (marc_d):

\[f(-2)=\frac{3}{-1}\times\frac{8}{8}\]

MARC (marc_d):

\[f(-2)=\frac{3}{-1}\times1\]

MARC (marc_d):

\[f(-2)=?\]solve it! @zachzach

OpenStudy (wanell):



OpenStudy (wanell):

plug -2 into X

OpenStudy (wanell):